a boy drags a 10kg box across a frictionless surface. the boy applies a force of 100N at an angle of 30° to the horizontal for 3 m in 5 seconds .
1.)calculate the work done by gravity
2. what power does the boy develop
1
Expert's answer
2020-06-29T14:03:45-0400

1)


"W=mgd\cos{30}\\W=(10)(9.8)(3)\cos{30}=255\ J"

2)


"P=\frac{Fd}{t}=\frac{(100)(3)}{5}=60\ W"